Weekly Inspection 10th August


Weather: improved and temperatures, 18° to 22°C. Some heavy showers,

Forage: golden rod, knot weed, rosebay willow herb and borage

Starting to see foraging activity slow down. lots of wasps around

Hive Record
CommentNext Inspection
AquaNYFour frames of bias, less Brood than previous inspectionLight only about 4 kgColony looks to be shrinking, not sure about QueenCheck on whether the colony is viable
BlueYYSix frames of biasOnly about 1 kg in second super, 1st super about 3/4 fullLight on stores in brood box, will need feedingCheck on feed required
CyanYYFive frames of biasOkay for stores, around 6-7 kgWill need feedingCheck on feed required
DenimNY7 1/2 frames of BIAS, plenty of eggsSecond super half fullLight on stores in brood box, only 4 to 5 kgCheck on feed required
EmeraldYY8+5 frames of bias, two brood frames replaced after donation to nuc4th super half fullAbout 8 kg in brood boxChecked on feed required
FlaxYYFive frames of bias, nice patternSecond, super 2/3 fullMarked QueenCheck on Queen
GreenYYFive frames of bias, starting to draw new framesLight on storesWill need feeding, check if continuing to draw framesCheck on frames being drawn
IndigoNYseven frames of bias with Good patternSuper 3/4 full15 kg in brood box, won’t need feedingCheck the Queen as not seen for a few weeks
JadeNYThree frames of BIAS, lots of eggsOkay for storesStarted drawing fifth frame Check on feed requirements